Circular No. 7411

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

CI AQUILAE = PROBABLE NOVA IN AQUILA H. Yamaoka, Kyushu University, reports an accurate position for the

object reported on IAUC 7409, measured from a V -band CCD image taken with Bisei Astronomical Observatory 1.01-m telescope by K. Ayani and K. Shirakami: α = 18^{h}52^{m}03^{s}.55, δ = −1^{o}28′38′′.9 (equinox 2000.0; uncertainty ∼ 0′′.2). This position is nearly identical to that of CI Aql.

W. Liller, Viña del Mar, Chile, reports that a prediscovery hypersensitized Technical Pan film, taken on Apr. 28.260 UT with an 85-mm f/2.0 camera lens and orange filter, clearly shows CI Aql at mag 9.8. He adds: “CCD spectroscopy with an objective prism shows Hα with an intensity 4.3 times that of the neighboring continuum. The [O I] lines at 630 and 636 nm are also strong, and there is evidence of He II at 587 nm. Hβ is present but weak, probably owing to interstellar reddening. There is no evidence (yet) of Fe II lines. CCD imaging with a 0.20-m Schmidt camera and a broadband-V filter (that includes the strong Hα emission) yields the following magnitudes: Apr. 30.360, 8.87; 30.398, 8.95.”

CCD photometry of CI Aql by D. Hanžl, Brno, Czech Republic, with a 0.40-m reflector: May 1.044 UT, V = 8.83 ± 0.03, B − V = +0.67 ± 0.03, V − I = +1.68 ± 0.05 (Hipparcos comparison star HIP 92546). Visual magnitude estimates: Apr. 30.063 UT, 9.4 (M. Reszelski, Szamotuly, Poland); 30.980, 9.1 (K. Hornoch, Lelekovice, Czech Republic); May 1.083, 8.9 (M. Lehký, Hradec Králové, Czech Republic).

COMET C/2000 G2 (LINEAR) Another object with unusual motion that was reported as asteroidal by

LINEAR (discovery observation given below) has been noted by other observers, following posting on the NEO Confirmation Page, as being cometary in appearance. G. Hug, Eskridge, KS, reported a diffuse appearance on several CCD images taken with a 0.3-m reflector during Apr. 21–29, and Kleť CCD observations by J. Tichá, M. Tichý, and Z. Moravec (0.57-m reflector) indicated a coma diameter of ∼ 6′′ on Apr. 22.9 UT. A 300-s R exposure taken by C. Hergenrother with the Steward Observatory 2.3-m telescope on Apr. 30 confirms that this object has a 6′′ coma and a 20′′ tail in p.a. 117^{o}. Full astrometry and orbital elements (T = 2000 Feb. 6.1 TT, q = 2.72 AU, i = 170^{o}.5, P = 53.6 yr) appear on MPEC 2000-J02.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Apr. 4.38585  14^{h}36^{m}04.^{s}21  −1^{o}57′13′′.1  18.4
2000 May 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green